# More with Less: An Empirical Study of Turn-Control Strategies for Efficient Coding Agents

### v1

@@ 1. Introduction
Large Language Models (LLMs) have profoundly impacted software development, with their most prominent application being the Coding Agent. A coding agent typically comprises an LLM at its core, a set of tools (e.g., a file editor, a shell), and an interactive environment that includes the project’s codebase. The agent operates in a multi-round loop to solve a given task, such as a GitHub issue. Each cycle in this loop, referred to as a turn, involves the LLM analyzing the problem state, selecting a tool to interact with the environment, executing it, and then processing the output to decide its next action. This loop continues until the agent determines the task is complete. This powerful paradigm is already deployed in popular AI-powered IDEs like Cursor (Cursor, 2025), TRAE (TRAE, 2025) and CLI assistants like Claude Code (Anthropic, 2025), and Gemini CLI (Google, 2025).
However, a critical challenge limiting the widespread, cost-effective deployment of these agents is their significant resource consumption, particularly in terms of tokens. We argue that managing the number of turns is a crucial, yet underexplored, strategy for controlling this cost. Our argument is built on the following observations: 1 First, the token cost of an agentic interaction grows quadratically with the number of turns. In every single typical agentic loop, the entire conversation history, including all previous prompts, tool calls, and their output, is fed back into the LLM as context for the next turn. This leads to a prompt size that grows non-linearly, often quadratically $(O(n^{2}))$, with the number of turns $n$, causing computational and financial costs to escalate rapidly. 2 Second, the state-of-the-art (SOTA) models best suited for these tasks are expensive. For instance, Anthropic’s Claude 4 Sonnet is priced at \$3.00 per million input tokens and \$15.00 per million output tokens. Its more powerful counterpart, Claude 4 Opus, costs \$15.00 and \$75.00 per million input and output tokens, respectively. When combined with the quadratic growth of tokens, high turn counts directly translate to prohibitive costs. 3 Third, solving real-world tasks requires a large number of turns. An analysis of top-performing agents on the SWE-bench Verified leaderboard (SWE-bench, 2025), a benchmark of real GitHub issues, illustrates this reality. As shown in Figure 1, the median number of turns required by five leading agents to solve an instance ranges from 41 to 58, with some complex tasks demanding over 175 turns. In our own experiments using Claude 4 Sonnet, an agent with no turn limit cost an average of \$5.85 to generate a patch and \$7.80 for a correct patch, underscoring the substantial expense. 4 Finally, agents are prone to wasting turns on unnecessary or even harmful actions. Without explicit constraints, models can be verbose or take indirect paths to a solution. For example, users on the Cursor IDE forum have reported agents adding "duplicate functions" and "redundant code" while fixing a simple bug, or breaking a project by attempting to refactor and add unrequested features (Cursor Community Forum, 2024). These behaviors not only inflate the turn count and cost but can also degrade the quality of the final output.
@@
While research into improving the efficiency of LLM-based agents is active, existing work has a unifying limitation: it focuses on optimizing the cost or content of individual turns, not the number of turns in the agentic loop. Current strategies include optimizing the prompt via compression (Jiang et al., 2023) and RAG (Lewis et al., 2020); re-architecting the agent with plan caching (Zhang et al., 2025) or multi-agent systems (Gandhi et al., 2024); and constraining the output with per-turn token budgets (Han et al., 2024). These techniques address the payload of each step, not the length of the entire workflow. The common use of a static step limit as a safeguard (Jimenez et al., 2023; Cognition AI, 2024) is a brute-force approach that lacks the adaptability needed for tasks of varying complexity. Consequently, the explicit study of dynamic turn-control strategies as a mechanism for cost-performance management in coding agents is not yet well-established.
To bridge this gap, we conduct a comprehensive empirical study to investigate the impact of various turn-control strategies on the performance and cost of coding agents. Our study is grounded in the realistic and challenging SWE-bench Verified benchmark, utilizing a representative subset of 100 tasks involving both bug fixes and feature implementations. We evaluate three SOTA closed-source models: Claude 4 Sonnet, Gemini 2.5 Pro, and GPT 4.1. We first establish a baseline with an infinite-turn agent. We then investigate a fixed-turn strategy, where agents operate under a hard deadline with explicit reminders of their remaining turns. Finally, we propose and evaluate a dynamic-turn strategy, which starts with a conservative turn budget and grants a one-time extension only when necessary. In addition to these quantitative experiments, we present in-depth case studies to qualitatively analyze the underlying behavioral changes that explain why constrained agents can outperform their unrestricted counterparts.
@@
RQ1: What are the performance and cost of coding agents in an unrestricted, infinite-turn setting? We first explore the baseline performance and cost of agents operating without any turn limitations, where the only exit condition is the agent’s own decision to terminate. This provides a crucial reference point for different LLMs and a baseline for our subsequent experiments.
@@ 2.4. Agent Design
We adopt TRAE Agent [^1] https://github.com/bytedance/trae-agent, the open-sourced and state-of-the-art agent on the SWE-bench Verified benchmark for our experiment. The agent is equipped with a specific toolset and can be deployed in one of three configurations.
@@
  • Infinite-Turn Agent (Baseline): This agent serves as our experimental baseline. Its inputs are the issue description and the project codebase. The agent operates in a continuous loop: the LLM generates a tool call. If the tool is task_done, the process terminates, and the patch is saved. Otherwise, the specified tool is executed, and its output is fed back to the LLM for the next turn. This agent has no turn restrictions.
@@ 3. Study Results
3.1. RQ1: Performance and Cost on Infinite-Turn Agent
@@
Table 2. Agent Performance Metrics in the Infinite-Turn Configuration
To answer RQ1, we conducted experiments using the infinite-turn agent configuration. The comprehensive results are presented in Table 2. The first column denotes the Large Language Model (LLM) used. The second column indicates the percentage of problems successfully solved by the agent. #Empty shows the number of instances where the agent failed to generate a patch. #Total Turns represents the cumulative number of conversational turns across all problems. The subsequent three columns, #Turn (25th), #Turn (50th), and #Turn (75th), detail the percentile distribution for the number of turns required to solve a single problem, with the median (50th) indicating typical performance. The #Input Tokens and #Output Tokens columns list the aggregate tokens consumed and generated, respectively. Finally, Total Cost presents the total economic expenditure in USD based on API pricing. Our analysis is structured around three core dimensions: problem-solving effectiveness, turn efficiency, and economic cost.
@@ 3.1.3. Economic Cost
Takeaways for RQ1 No One-Size-Fits-All Model: There is no single best model; selection depends on the specific priorities of the task, balancing effectiveness, efficiency, and cost. For Maximum Effectiveness: Claude 4 Sonnet is the top performer, achieving the highest solve rate (75%). This makes it the ideal choice when the primary goal is to maximize the success of problem-solving, despite its higher cost. For Optimal Cost-Efficiency: GPT 4.1 offers the best value. It is the most budget-friendly option and delivers a solid solve rate (62%), matching Gemini 2.5 Pro’s performance at a significantly lower cost.
@@ 3.2. RQ2: Performance and Cost on Fixed-Turn Agent with Reminder
To address RQ2, we investigate how imposing a fixed turn limit, coupled with a number of turns left reminder fed to models, affects model performance and cost. We constrained the agents to turn limits corresponding to the 25th, 50th, and 75th percentiles of the turns consumed by their respective baseline runs (c.f. Table 2). The detailed results of this experiment are presented in Table 3, which contrasts the performance of the constrained agents against the unconstrained baseline. The table is structured into four sections for clear comparison. The first section (Rows 3-5) establishes the baseline performance from the unconstrained, infinite-turn configuration discussed in RQ1. The subsequent sections display the results when the turn limit is fixed at the 25th (Rows 7-9), 50th (Rows 11-13), and 75th (Rows 15-17) percentiles of each model’s baseline turn consumption. Within the table, the first column denotes the LLM. Solve Rate is the percentage of successfully solved problems. #Empty counts failed attempts that yielded an empty patch. #Total Turns, #Input Tokens, and #Output Tokens quantify resource consumption, and Total Cost shows the final expenditure in USD. Crucially, the values in parentheses () in the Solve Rate and Total Cost columns indicate the percentage change relative to the infinite-turn baseline, making it easy to assess performance impact and cost reduction.
@@ 3.2.3. Synthesized Analysis: Identifying the Optimal Efficacy-Cost Ratio
Takeaways for RQ2 Primarily a Cost-Optimization Tool: The fixed-turn strategy’s main value lies in its ability to enforce budgetary control. It consistently yields significant cost savings, typically in exchange for a controlled degradation in solve rate. Optimal "Sweet Spot" at 75th Percentile: Across all models, setting the turn limit to the 75th percentile of the baseline distribution provides the most favorable efficacy-cost ratio. This setting secures substantial cost reductions (ranging from 23% to 68%) while incurring only minimal performance losses. The "More with Less" Anomaly: The combination of Gemini 2.5 Pro and a 75th-percentile limit is a standout exception. It not only cuts costs by over 68% but also improves the solve rate by 3.17%, demonstrating that a moderate sense of urgency can catalyze a highly efficient model to unlock its potential. Actionable Recommendation: For deploying agentic systems in budget-constrained environments, we strongly recommend implementing a fixed-turn limit. A setting at approximately the 75th percentile of the expected turn distribution is a robust strategy to dramatically reduce operational costs with negligible, and sometimes even positive, impact on performance.
@@ 3.3.3. Discussion: The Principle of Efficient Resource Allocation
Takeaways for RQ3 Unambiguously Superior Strategy: The dynamic-turn strategy consistently outperforms its fixed-turn counterparts in all tested scenarios. It is demonstrably more efficient and effective. Achieves More with Less: This approach provides a reliable method to significantly reduce costs (by 5% to 24%) while simultaneously preserving—and in some cases, even improving—the problem-solving success rate. Based on Efficient Resource Allocation: Its success stems from allocating resources only when needed. It avoids the waste inherent in fixed-limit strategies by allowing easy tasks to finish cheaply and providing a "second chance" only to tasks that genuinely require it. Recommended for Practical Deployment: For any agentic system where balancing high performance with cost management is critical, a dynamic, multi-stage turn allocation should be the default and preferred implementation strategy.
@@ 4.1. Case Study 1
Our quantitative results show that Gemini 2.5 Pro under a 75th-percentile fixed-turn limit with reminders, achieved a 65% solve rate, surpassing the 63% solve rate of its unrestricted, infinite-turn counterpart. To understand the underlying behavioral changes driving this performance improvement, we present a case study of a instanced named sympy__sympy-22456. This instance is particularly illustrative: the 75th-percentile agent successfully generated a correct patch in just 35 turns, whereas the infinite-turn agent failed after a prolonged 82-turn trajectory that produced an incorrect patch. Analysis of these two trajectories reveals how moderate pressure can guide an agent toward a more effective problem-solving path.
@@
In contrast, the infinite-turn agent’s trajectory exemplifies a cascading failure initiated by a flawed diagnosis. The critical divergence occurred at the initial bug-reproduction step. The agent received a more ambiguous error message, TypeError: No value for ’text’ given and attribute has no default, and observed that s.args was an empty tuple. Instead of investigating why the arguments were empty, it incorrectly focused on the TypeError as the primary symptom, hypothesizing that the issue lay in how s.func() was called with no arguments. This misdiagnosis led it down an unproductive path of erroneous fixes. It spent dozens of turns repeatedly rewriting the __new__ method in both the String class and its parents, introducing new NameError and AttributeError exceptions. This created a vicious cycle where the agent was forced to address self-inflicted errors, such as fixing a TypeError that led to a new problem with __eq__, which in turn created an issue with __hash__. Having lost the original objective, the agent even attempted to modify unrelated files like cnodes.py and became trapped in loops of undoing and retrying its own faulty modifications. The agent’s reasoning became polluted by a history of failed attempts and contradictory hypotheses, making it increasingly difficult to return to a correct line of reasoning.

### v2

@@ Abstract.
Bug Fixing, Coding Agents, Large Language Models, Automated Program Repair
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773250isbn: 979-8-4007-2025-3/26/04ccs: Software and its engineering Software verification and validation
@@ 1. Introduction
Large Language Models (LLMs) have profoundly impacted software development, with their most prominent application being the Coding Agent. A coding agent typically comprises an LLM at its core, a set of tools (e.g., a file editor, a shell), and an interactive environment that includes the project’s codebase. The agent operates in a multi-round loop to solve a given task, such as a GitHub issue. Each cycle in this loop, referred to as a turn, involves the LLM analyzing the problem state, selecting a tool to interact with the environment, executing it, and then processing the output to decide its next action. This loop continues until the agent determines the task is complete. This powerful paradigm is already deployed in popular AI-powered IDEs like Cursor (cursor), TRAE (traeagent2025) and CLI assistants like Claude Code (claude_code), and Gemini CLI (gemini_cli).
@@
However, a critical challenge limiting the widespread, cost-effective deployment of these agents is their significant resource consumption, particularly in terms of tokens. We argue that managing the number of turns is a crucial, yet underexplored, strategy for controlling this cost. Our argument is built on the following observations: 1 First, the token cost of an agentic interaction can grow faster than linearly with the number of turns. In every single typical agentic loop, the entire conversation history, including all previous prompts, tool calls, and their output, is fed back into the LLM as context for the next turn. This leads to a prompt size that grows non-linearly, often quadratically $(O(n^{2}))$, with the number of turns $n$, causing computational and financial costs to escalate rapidly. 2 Second, the state-of-the-art (SOTA) models best suited for these tasks are expensive. For instance, Anthropic’s Claude 4 Sonnet is priced at \$3.00 per million input tokens and \$15.00 per million output tokens. Its more powerful counterpart, Claude 4 Opus, costs \$15.00 and \$75.00 per million input and output tokens, respectively. When combined with the quadratic growth of tokens, high turn counts directly translate to prohibitive costs. 3 Third, solving real-world tasks requires a large number of turns. An analysis of top-performing agents on the SWE-bench Verified leaderboard (swebenchleaderboard), a benchmark of real GitHub issues, illustrates this reality. As shown in Figure 1, the median number of turns required by five leading agents to solve an instance ranges from 41 to 58, with some complex tasks demanding over 175 turns. In our own experiments using Claude 4 Sonnet, an agent with no turn limit cost an average of \$5.85 to generate a patch and \$7.80 for a correct patch, underscoring the substantial expense. 4 Finally, agents are prone to wasting turns on unnecessary or even harmful actions. Without explicit constraints, models can be verbose or take indirect paths to a solution. For example, users on the Cursor IDE forum have reported agents adding "duplicate functions" and "redundant code" while fixing a simple bug, or breaking a project by attempting to refactor and add unrequested features (cursorforum2024). These behaviors not only inflate the turn count and cost but can also degrade the quality of the final output.
@@
While research into improving the efficiency of LLM-based agents is active, existing work has a unifying limitation: it focuses on optimizing the cost or content of individual turns, not the number of turns in the agentic loop. Current strategies include optimizing the prompt via compression (jiang2023llmlingua) and RAG (lewis2020rag); re-architecting the agent with plan caching (zhang2025cost) or multi-agent systems (gandhi2024budgetmlagent); and constraining the output with per-turn token budgets (han2024token). These techniques address the payload of each step, not the length of the entire workflow. The common use of a static step limit as a safeguard (jimenez2023swebench; cognition2024swebench) is a brute-force approach that lacks the adaptability needed for tasks of varying complexity. Consequently, the explicit study of dynamic turn-control strategies as a mechanism for cost-performance management in coding agents is not yet well-established.
To bridge this gap, we conduct a comprehensive empirical study to investigate the impact of various turn-control strategies on the performance and cost of coding agents. Our study is grounded in the realistic and challenging SWE-bench Verified benchmark, utilizing a representative subset of 100 tasks involving both bug fixes and feature implementations. We evaluate three SOTA closed-source models: Claude 4 Sonnet, Gemini 2.5 Pro, and GPT 4.1. We first establish a baseline with an unlimited-turn agent. We then investigate a fixed-turn strategy, where agents operate under a hard deadline with explicit reminders of their remaining turns. Finally, we propose and evaluate a dynamic-turn strategy, which starts with a conservative turn budget and grants a one-time extension only when necessary. In addition to these quantitative experiments, we present in-depth case studies to qualitatively analyze the underlying behavioral changes that explain why constrained agents can outperform their unrestricted counterparts.
@@
RQ1: What are the performance and cost of coding agents in an unrestricted, unlimited-turn setting? We first explore the baseline performance and cost of agents operating without any turn limitations, where the only exit condition is the agent’s own decision to terminate. This provides a crucial reference point for different LLMs and a baseline for our subsequent experiments.
@@ 2.4. Agent Design
We adopt TRAE Agent(traeagent2025) [^1] https://github.com/bytedance/trae-agent, the open-sourced and state-of-the-art agent on the SWE-bench Verified benchmark for our experiment. The agent is equipped with a specific toolset and can be deployed in one of three configurations.
@@
  • Unlimited-Turn Agent (Baseline): This agent serves as our experimental baseline. Its inputs are the issue description and the project codebase. The agent operates in a continuous loop: the LLM generates a tool call. If the tool is task_done, the process terminates, and the patch is saved. Otherwise, the specified tool is executed, and its output is fed back to the LLM for the next turn. This agent has no turn restrictions.
@@ 3. Study Results
3.1. RQ1: Performance and Cost on Unlimited-Turn Agent
Table 2. Agent Performance Metrics in the Unlimited-Turn Configuration
@@ 3.1. RQ1: Performance and Cost on Unlimited-Turn Agent
To answer RQ1, we conducted experiments using the unlimited-turn agent configuration. The comprehensive results are presented in Table 2. The first column denotes the Large Language Model (LLM) used. The second column indicates the percentage of problems successfully solved by the agent. #Empty shows the number of instances where the agent failed to generate a patch. #Total Turns represents the cumulative number of conversational turns across all problems. The subsequent three columns, #Turn (25th), #Turn (50th), and #Turn (75th), detail the percentile distribution for the number of turns required to solve a single problem, with the median (50th) indicating typical performance. The #Input Tokens and #Output Tokens columns list the aggregate tokens consumed and generated, respectively. Finally, Total Cost presents the total economic expenditure in USD based on API pricing. Our analysis is structured around three core dimensions: problem-solving effectiveness, turn efficiency, and economic cost.
@@ 3.1.3. Economic Cost
Takeaway for RQ1
  • No One-Size-Fits-All Model: There is no single best model; selection depends on the specific priorities of the task, balancing effectiveness, efficiency, and cost.
  • For Maximum Effectiveness: Claude 4 Sonnet is the top performer, achieving the highest solve rate (75%). This makes it the ideal choice when the primary goal is to maximize the success of problem-solving, despite its higher cost.
  • For Optimal Cost-Efficiency: GPT 4.1 offers the best value. It is the most budget-friendly option and delivers a solid solve rate (62%), matching Gemini 2.5 Pro’s performance at a significantly lower cost.
@@ 3.2. RQ2: Performance and Cost on Fixed-Turn Agent with Reminder
To address RQ2, we investigate how imposing a fixed turn limit, coupled with a number of turns left reminder fed to models, affects model performance and cost. We constrained the agents to turn limits corresponding to the 25th, 50th, and 75th percentiles of the turns consumed by their respective baseline runs (c.f. Table 2). The detailed results of this experiment are presented in Table 3, which contrasts the performance of the constrained agents against the unconstrained baseline. The table is structured into four sections for clear comparison. The first section (Rows 3-5) establishes the baseline performance from the unconstrained, unlimited-turn configuration discussed in RQ1. The subsequent sections display the results when the turn limit is fixed at the 25th (Rows 7-9), 50th (Rows 11-13), and 75th (Rows 15-17) percentiles of each model’s baseline turn consumption. Within the table, the first column denotes the LLM. Solve Rate is the percentage of successfully solved problems. #Empty counts failed attempts that yielded an empty patch. #Total Turns, #Input Tokens, and #Output Tokens quantify resource consumption, and Total Cost shows the final expenditure in USD. Crucially, the values in parentheses () in the Solve Rate and Total Cost columns indicate the percentage change relative to the unlimited-turn baseline, making it easy to assess performance impact and cost reduction.
@@ 3.2.1. Performance Impact: Convergence with a Cost
3.2.2. Agent Performance without Turn Reminders
In a scenario where the agent is not reminded of the turn limit, it would simply halt upon exhausting its turns, resulting in an empty patch and a failed task. Therefore, we re-analyzed our baseline data: if an instance was successfully solved but required more turns than the given limit, we re-classified it as a failure with an empty patch. The results, summarized in Table 4, show a clear and significant performance degradation (less solve rate and more empty patches) when no reminders are given.
Table 4. Agent Performance with and without Turn Reminders
Percentile LLM Solve Rate (w/ rem.) Solve Rate (w/o rem.) #Empty (w/ rem.) #Empty (w/o rem.)
25th Claude 4 Sonnet 60% 24% (↓ 60%) 16 73 (↑ 57)
Gemini 2.5 Pro 26% 20% (↓ 23%) 63 73 (↑ 10)
GPT 4.1 49% 24% (↓ 51%) 26 71 (↑ 45)
50th Claude 4 Sonnet 63% 46% (↓ 27%) 11 48 (↑ 37)
Gemini 2.5 Pro 52% 39% (↓ 25%) 24 49 (↑ 25)
GPT 4.1 57% 40% (↓ 30%) 10 48 (↑ 38)
75th Claude 4 Sonnet 71% 63% (↓ 11%) 3 25 (↑ 22)
Gemini 2.5 Pro 65% 54% (↓ 17%) 4 25 (↑ 21)
GPT 4.1 57% 54% (↓ 5%) 3 24 (↑ 21)
Note: Rem. is the abbreviation of reminder.
@@ 3.2.4. Synthesized Analysis: Identifying the Optimal Efficacy-Cost Ratio
Takeaway for RQ2
  • Primarily a Cost-Optimization Tool: The fixed-turn strategy’s main value lies in its ability to enforce budgetary control. It consistently yields significant cost savings, typically in exchange for a controlled degradation in solve rate.
  • Optimal "Sweet Spot" at 75th Percentile: Across all models, setting the turn limit to the 75th percentile of the baseline distribution provides the most favorable efficacy-cost ratio. This setting secures substantial cost reductions (ranging from 23% to 68%) while incurring only minimal performance losses.
  • The "More with Less" Anomaly: The combination of Gemini 2.5 Pro and a 75th-percentile limit is a standout exception. It not only cuts costs by over 68% but also improves the solve rate by 3.17%, demonstrating that a moderate sense of urgency can catalyze a highly efficient model to unlock its potential.
  • Actionable Recommendation: For deploying agentic systems in budget-constrained environments, we strongly recommend implementing a fixed-turn limit. A setting at approximately the 75th percentile of the expected turn distribution is a robust strategy to dramatically reduce operational costs with negligible, and sometimes even positive, impact on performance.
@@ 3.3.3. Discussion: The Principle of Efficient Resource Allocation
Takeaway for RQ3
  • Unambiguously Superior Strategy: The dynamic-turn strategy consistently outperforms its fixed-turn counterparts in all tested scenarios. It is demonstrably more efficient and effective.
  • Achieves More with Less: This approach provides a reliable method to significantly reduce costs (by 5% to 24%) while simultaneously preserving—and in some cases, even improving—the problem-solving success rate.
  • Based on Efficient Resource Allocation: Its success stems from allocating resources only when needed. It avoids the waste inherent in fixed-limit strategies by allowing easy tasks to finish cheaply and providing a "second chance" only to tasks that genuinely require it.
  • Recommended for Practical Deployment: For any agentic system where balancing high performance with cost management is critical, a dynamic, multi-stage turn allocation should be the default and preferred implementation strategy.
@@ 4.1. Case Study 1
Our quantitative results show that Gemini 2.5 Pro under a 75th-percentile fixed-turn limit with reminders, achieved a 65% solve rate, surpassing the 63% solve rate of its unrestricted, unlimited-turn counterpart. To understand the underlying behavioral changes driving this performance improvement, we present a case study of a instanced named sympy__sympy-22456. This instance is particularly illustrative: the 75th-percentile agent successfully generated a correct patch in just 35 turns, whereas the unlimited-turn agent failed after a prolonged 82-turn trajectory that produced an incorrect patch. Analysis of these two trajectories reveals how moderate pressure can guide an agent toward a more effective problem-solving path.
@@
In contrast, the unlimited-turn agent’s trajectory exemplifies a cascading failure initiated by a flawed diagnosis. The critical divergence occurred at the initial bug-reproduction step. The agent received a more ambiguous error message, TypeError: No value for ’text’ given and attribute has no default, and observed that s.args was an empty tuple. Instead of investigating why the arguments were empty, it incorrectly focused on the TypeError as the primary symptom, hypothesizing that the issue lay in how s.func() was called with no arguments. This misdiagnosis led it down an unproductive path of erroneous fixes. It spent dozens of turns repeatedly rewriting the __new__ method in both the String class and its parents, introducing new NameError and AttributeError exceptions. This created a vicious cycle where the agent was forced to address self-inflicted errors, such as fixing a TypeError that led to a new problem with __eq__, which in turn created an issue with __hash__. Having lost the original objective, the agent even attempted to modify unrelated files like cnodes.py and became trapped in loops of undoing and retrying its own faulty modifications. The agent’s reasoning became polluted by a history of failed attempts and contradictory hypotheses, making it increasingly difficult to return to a correct line of reasoning.